OpenStudy (anonymous):

Factor the following trinomial completely: -3x^3 + 66x^2 - 363x A. -3x(x - 11)(x + 11) B. -3x(x - 11)^2 C. 3x(x - 11)(x + 11) D. 3x(x - 11)^2

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (jdoe0001):

hint: \(\large { -3x^3 + 66x^2 - 363x\implies \begin{array}{cccllll} -3x(x^2&-22x&+121)\\ &-11-11&-11\cdot -11 \end{array} }\)

OpenStudy (anonymous):

Okay so what should I do with that next

OpenStudy (michele_laino):

hint: the polynomial: \[{x^2} - 22x + 121\] is a perfect square

OpenStudy (michele_laino):

for example: \[\Large {\left( {x - a} \right)^2} = {x^2} - 22x + 121\] what is a?

OpenStudy (anonymous):

How would I find that?

OpenStudy (michele_laino):

hint: \[\Large {\left( {x - a} \right)^2} = {x^2} - 2ax + {a^2}\]

OpenStudy (michele_laino):

so we can write: \[\Large {x^2} - 2ax + {a^2} = {x^2} - 22x + 121\]

OpenStudy (michele_laino):

using the principle of identity of polynomial, we can write: \[\Large \begin{gathered} - 2a = - 22 \hfill \\ {a^2} = 121 \hfill \\ \end{gathered} \] so, what is a?

OpenStudy (anonymous):

11!

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so the requested factorization, is: \[ \Large - 3x\left( {{x^2} - 22x + 121} \right) = - 3x{\left( {x - 11} \right)^2}\]

OpenStudy (michele_laino):

is it ok?

OpenStudy (anonymous):

Yes!

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

So it's B

OpenStudy (michele_laino):

That's right!

OpenStudy (anonymous):

Awesome thanks! could you help with another problem? I'll put it in a new question and tag you

OpenStudy (michele_laino):

ok!